# Physics-Informed Teleconnection-Aware Transformer for Global Subseasonal-to-Seasonal Forecasting

### v2

@@ Appendix A Theoretical Analysis
A.1 Spectral Representation on the Sphere
The Spherical Harmonic Embedding (SHE) in our model builds upon the well-established spectral theory for functions on the sphere. Any square-integrable function $\color[rgb]{0.171875,0.2421875,0.3125}f(\theta,\phi)$ on the sphere can be represented as an infinite series of spherical harmonics:
$f(\theta,\phi)=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}a_{\ell m}Y_{\ell}^{m}(\theta,\phi)$ (10)
where $\color[rgb]{0.171875,0.2421875,0.3125}Y_{\ell}^{m}(\theta,\phi)$ are the spherical harmonics of degree $\color[rgb]{0.171875,0.2421875,0.3125}\ell$ and order $\color[rgb]{0.171875,0.2421875,0.3125}m$, and $\color[rgb]{0.171875,0.2421875,0.3125}a_{\ell m}$ are the spectral coefficients. Our learnable zonal averaging approach in SHE provides a computationally efficient approximation to the important subset of zonal harmonics ($\color[rgb]{0.171875,0.2421875,0.3125}m=0$). This connection is justified by the observations in atmospheric science that many large-scale circulation patterns exhibit strong zonal symmetry baldwin2003stratospheric ; kidston2015stratospheric .
Proposition 1 (Zonal Projection Property).
[…]
$\bar{f}(\theta)=\sum_{\ell=0}^{\infty}a_{\ell,0}Y_{\ell}^{0}(\theta)$ (11)
where $\color[rgb]{0.171875,0.2421875,0.3125}Y_{\ell}^{0}(\theta)=\sqrt{\frac{2\ell+1}{4\pi}}P_{\ell}(\cos\theta)$ are the zonal harmonics and $\color[rgb]{0.171875,0.2421875,0.3125}P_{\ell}$ are the Legendre polynomials.
Proposition 2 (Approximation Quality).
@@ A.1 Spectral Representation on the Sphere
$\left\|\mathbf{u}_{i}-\sum_{\ell=0}^{L}a_{\ell,0}P_{\ell}(\cos\theta_{i})\right\|\leq\mathcal{O}(D^{-1/2}+L^{-1})$ (12)
where $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{u}_{i}=\frac{1}{W}\sum_{j=1}^{W}\mathbf{X}_{i,j}$ is the empirical zonal average and the approximation improves with embedding dimension $\color[rgb]{0.171875,0.2421875,0.3125}D$ and truncation level $\color[rgb]{0.171875,0.2421875,0.3125}L$.
@@
A.2 Multi-Scale Analysis and Function Approximation
@@
Assumption 1 (Signal Regularity). We assume that atmospheric signals possess sufficient regularity, specifically Hölder continuity with exponent $\color[rgb]{0.171875,0.2421875,0.3125}\alpha>0$, enabling the stated approximation rates. This assumption is well-justified for large-scale atmospheric variables due to the underlying physics.
Proposition 3 (Task-Adaptive Decomposition).
@@
$\mathbb{E}[\|\mathbf{x}-\sum_{\ell=0}^{L}\mathbf{x}^{(\ell)}\|^{2}]\leq C\cdot L^{-\alpha}$ (13)
where $\color[rgb]{0.171875,0.2421875,0.3125}C$ is a constant depending on the signal characteristics and $\color[rgb]{0.171875,0.2421875,0.3125}\alpha>0$ reflects the smoothness of atmospheric signals. $\color[rgb]{0.171875,0.2421875,0.3125}L$ is the number of frequency bands.
@@
The MLP transforms at each level provide universal function approximation capabilities with sufficient width. By minimizing the end-to-end prediction loss, these MLPs learn to extract frequency components that maximize predictive skill rather than reconstruction fidelity. The recursive binary splitting operation ensures that the $\color[rgb]{0.171875,0.2421875,0.3125}L+1$ frequency bands form a partition of the signal’s spectral content. Under the regularity assumption, this adaptive partitioning achieves the stated convergence rate through standard approximation theory arguments. ∎
Proposition 4 (Information Preservation).
Each frequency band $\color[rgb]{0.171875,0.2421875,0.3125}\ell$ preserves mutual information with the forecast target, satisfying:
$\sum_{\ell=0}^{L}I(\mathbf{x}^{(\ell)};\mathbf{y})\geq(1-\delta)I(\mathbf{x};\mathbf{y})$ (14)
for arbitrarily small $\color[rgb]{0.171875,0.2421875,0.3125}\delta>0$ given sufficient model capacity, where $\color[rgb]{0.171875,0.2421875,0.3125}I(\cdot;\cdot)$ denotes mutual information.
@@
A.3 Physics-Informed Neural ODEs
@@
Latent Space Transport Dynamics: We model the evolution of latent representations $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{x}_{i}\in\mathbb{R}^{D}$ along the latitude dimension using a transport-inspired ODE:
$\frac{d\mathbf{x}_{i}}{dt}=\gamma\cdot\tanh\Bigg{(}{\color[rgb]{0.0078125,0.46484375,0.7421875}\underbrace{\boldsymbol{\nu}\odot(\mathbf{x}_{i+1}-2\mathbf{x}_{i}+\mathbf{x}_{i-1})}_{\smash{\text{diffusion}}}}+{\color[rgb]{0.84765625,0.26171875,0.08203125}\underbrace{\boldsymbol{\mu}\odot\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i-1}}{2}}_{\smash{\text{advection}}}}+{\color[rgb]{0.1875,0.24609375,0.625}\underbrace{\boldsymbol{f}}_{\smash{\text{forcing}}}}+{\color[rgb]{0,0.5390625,0.48046875}\underbrace{\alpha\cdot\text{MLP}(\mathbf{x}_{i})}_{\smash{\text{neural correction}}}}\Bigg{)}$ (15)
Proposition 5 (Latent Space Transport Interpretation).
@@
$\frac{d\mathbf{x}}{dt}=\mathcal{T}_{transport}(\mathbf{x};\boldsymbol{\theta}_{phys})+\mathcal{T}_{neural}(\mathbf{x};\boldsymbol{\theta}_{nn})$ (16)
where $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{T}_{transport}$ captures simplified transport processes with learnable physics parameters $\color[rgb]{0.171875,0.2421875,0.3125}\boldsymbol{\theta}_{phys}=\{\boldsymbol{\nu},\boldsymbol{\mu},\boldsymbol{f}\}\in\mathbb{R}^{3D}$, and $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{T}_{neural}$ provides additional modeling capacity through neural network parameters $\color[rgb]{0.171875,0.2421875,0.3125}\boldsymbol{\theta}_{nn}$.
@@
[…]
Proposition 6 (Numerical Stability).
The transport-inspired ODE with hyperbolic tangent activation is globally Lipschitz continuous with constant $\color[rgb]{0.171875,0.2421875,0.3125}L=\gamma$, ensuring controlled error propagation:
$\left\|\frac{d\mathbf{x}}{dt}-\frac{d\mathbf{x}^{\prime}}{dt}\right\|\leq\gamma\|\mathbf{x}-\mathbf{x}^{\prime}\|$ (18)
for any latent states $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{HD}$. This global Lipschitz property guarantees existence and uniqueness of solutions and supports numerical stability during integration.
Proposition 7 (Parameter Efficiency).
The physics-inspired structure reduces the effective parameter space by constraining the dynamics to transport-consistent manifolds. Specifically, the constraint reduces the degrees of freedom from $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{O}(HD^{2})$ (unconstrained) to $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{O}(HD+D^{2})$ (transport-constrained), potentially improving sample efficiency.
Implementation Considerations: We apply periodic boundary conditions consistent with our latitude-based processing and employ adaptive step-size integration methods for numerical accuracy. The scaling factor $\color[rgb]{0.171875,0.2421875,0.3125}\gamma$ and hyperbolic tangent activation prevent gradient explosion while preserving the fundamental transport structure.
@@
A.4 Teleconnection Representation Theory
@@
Proposition 8 (Nonlinear Teleconnection Learning).
The learned teleconnection patterns $\color[rgb]{0.171875,0.2421875,0.3125}\{\mathbf{P}_{j}\}_{j=1}^{n_{p}}\subset\mathbb{R}^{HD}$ converge to a basis that captures the dominant modes of atmospheric covariance, encompassing both linear relationships and nonlinear teleconnection patterns. Formally:
$\lim_{T\to\infty}\text{span}(\{\mathbf{P}_{j}\})\supseteq\text{span}(\{\text{EOF}_{j}^{(linear)}\})$ (19)
@@
The teleconnection attention computes pattern weights as $\color[rgb]{0.171875,0.2421875,0.3125}\boldsymbol{\omega}=\text{softmax}(\mathbf{\bar{x}}\mathbf{W}^{p})$ where $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{\bar{x}}\in\mathbb{R}^{D}$ is the global mean state. During training, the patterns $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{P}_{j}$ and projection matrix $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{W}^{p}\in\mathbb{R}^{D\times n_{p}}$ are jointly optimized to minimize prediction error.
@@
Proposition 9 (Sample Complexity for Teleconnection Learning).
To learn $\color[rgb]{0.171875,0.2421875,0.3125}k$ dominant teleconnection patterns with approximation error $\color[rgb]{0.171875,0.2421875,0.3125}\varepsilon$ in the $\color[rgb]{0.171875,0.2421875,0.3125}L_{2}$ norm, the teleconnection module requires $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{O}(k\log(D)/\varepsilon^{2})$ training samples, where the logarithmic dependence on dimension reflects the low-intrinsic-dimensional nature of atmospheric teleconnections.
Proposition 10 (Convergence Rate).
Under standard regularity conditions on the atmospheric data distribution, the teleconnection patterns converge to their population optima at rate $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{O}(T^{-1/2})$ where $\color[rgb]{0.171875,0.2421875,0.3125}T$ is the number of training iterations, matching the convergence rate of stochastic gradient methods for smooth objectives.
@@
A.5 Complexity and Generalization Analysis
@@
A.5.1 Computational Complexity
Proposition 11 (Time Complexity).
@@ A.5.1 Computational Complexity
A.5.2 Error Propagation Analysis
Proposition 12 (Multi-Scale Error Bounds).
For a signal decomposed into $\color[rgb]{0.171875,0.2421875,0.3125}L+1$ frequency bands, the prediction error is bounded by:
$\mathbb{E}[\|y-\hat{y}\|^{2}]\leq\sum_{\ell=0}^{L}\lambda_{\ell}\mathbb{E}[\|y_{\ell}-\hat{y}_{\ell}\|^{2}]+\mathcal{O}(L^{-\alpha})$ (21)
where $\color[rgb]{0.171875,0.2421875,0.3125}y_{\ell},\hat{y}_{\ell}$ represent ground truth and prediction for frequency band $\color[rgb]{0.171875,0.2421875,0.3125}\ell$, $\color[rgb]{0.171875,0.2421875,0.3125}\lambda_{\ell}\geq 0$ are band-specific weights with $\color[rgb]{0.171875,0.2421875,0.3125}\sum_{\ell}\lambda_{\ell}=1$, and the additional term bounds decomposition error.
@@ A.5.2 Error Propagation Analysis
The multi-scale architecture processes frequency bands independently before fusion, creating a natural error decomposition. The hierarchical fusion mechanism that privileges low-frequency information creates weights $\color[rgb]{0.171875,0.2421875,0.3125}\lambda_{\ell}$ that typically decrease with frequency, providing robustness against high-frequency noise. The decomposition error term reflects the finite precision of the learned wavelet basis. ∎
A.5.3 Generalization Bounds
Proposition 13 (Generalization Error).
For a training set of size $\color[rgb]{0.171875,0.2421875,0.3125}n$ drawn from the atmospheric state distribution, with probability at least $\color[rgb]{0.171875,0.2421875,0.3125}1-\delta$, the generalization error satisfies:
$\mathcal{R}(h)-\hat{\mathcal{R}}_{n}(h)\leq\mathcal{O}\left(\sqrt{\frac{\log(L+1)+\sum_{\ell=0}^{L}\mathcal{C}(h_{\ell})+\log(1/\delta)}{n}}\right)$ (22)
where $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{R}(h)$ is the true risk, ^$\color[rgb]{0.171875,0.2421875,0.3125}\hat{\mathcal{R}}_{n}(h)$ is the empirical risk, and $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{C}(h_{\ell})$ measures the complexity of the model component for frequency band $\color[rgb]{0.171875,0.2421875,0.3125}\ell$.

### v3

@@ Appendix A Theoretical Analysis
A.1. Spectral Representation on the Sphere
The Spherical Harmonic Embedding (SHE) in our model builds upon the well-established spectral theory for functions on the sphere. Any square-integrable function $f(\theta,\phi)$ on the sphere can be represented as an infinite series of spherical harmonics:
(21) $f(\theta,\phi)=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}a_{\ell m}Y_{\ell}^{m}(\theta,\phi)$
where $Y_{\ell}^{m}(\theta,\phi)$ are the spherical harmonics of degree $\ell$ and order $m$, and $a_{\ell m}$ are the spectral coefficients. Our learnable zonal averaging approach in SHE provides a computationally efficient approximation to the important subset of zonal harmonics ($m=0$). This connection is justified by the observations in atmospheric science that many large-scale circulation patterns exhibit strong zonal symmetry (Baldwin et al., 2003; Kidston et al., 2015).
Proposition A.1 (Zonal Projection Property).
The zonal averaging operation implements an orthogonal projection onto the subspace of zonally symmetric functions. For a function $f(\theta,\phi)$, the zonal average $\bar{f}(\theta)=\frac{1}{2\pi}\int_{-\pi}^{\pi}f(\theta,\phi)d\phi$ projects onto the $m=0$ spherical harmonic modes:
(22) $\bar{f}(\theta)=\sum_{\ell=0}^{\infty}a_{\ell,0}Y_{\ell}^{0}(\theta)$
where $Y_{\ell}^{0}(\theta)=\sqrt{\frac{2\ell+1}{4\pi}}P_{\ell}(\cos\theta)$ are the zonal harmonics and $P_{\ell}$ are the Legendre polynomials.
Proposition A.2 (Approximation Quality).
@@ A.1. Spectral Representation on the Sphere
(23) $\left\|\mathbf{u}_{i}-\sum_{\ell=0}^{L}a_{\ell,0}P_{\ell}(\cos\theta_{i})\right\|\leq\mathcal{O}(D^{-1/2}+L^{-1})$
where $\mathbf{u}_{i}=\frac{1}{W}\sum_{j=1}^{W}\mathbf{X}_{i,j}$ is the empirical zonal average and the approximation improves with embedding dimension $D$ and truncation level $L$.
@@
A.2. Multi-Scale Analysis and Function Approximation
@@
Assumption 1 (Signal Regularity). We assume that atmospheric signals possess sufficient regularity, specifically Hölder continuity with exponent $\alpha>0$, enabling the stated approximation rates. This assumption is well-justified for large-scale atmospheric variables due to the underlying physics.
Proposition A.3 (Task-Adaptive Decomposition).
@@
(24) $\mathbb{E}[\|\mathbf{x}-\sum_{\ell=0}^{L}\mathbf{x}^{(\ell)}\|^{2}]\leq C\cdot L^{-\alpha}$
where $C$ is a constant depending on the signal characteristics and $\alpha>0$ reflects the smoothness of atmospheric signals. $L$ is the number of frequency bands.
@@
The MLP transforms at each level provide universal function approximation capabilities with sufficient width. By minimizing the end-to-end prediction loss, these MLPs learn to extract frequency components that maximize predictive skill rather than reconstruction fidelity. The recursive binary splitting operation ensures that the $L+1$ frequency bands form a partition of the signal’s spectral content. Under the regularity assumption, this adaptive partitioning achieves the stated convergence rate through standard approximation theory arguments. ∎
Proposition A.4 (Information Preservation).
Each frequency band $\ell$ preserves mutual information with the forecast target, satisfying:
(25) $\sum_{\ell=0}^{L}I(\mathbf{x}^{(\ell)};\mathbf{y})\geq(1-\delta)I(\mathbf{x};\mathbf{y})$
for arbitrarily small $\delta>0$ given sufficient model capacity, where $I(\cdot;\cdot)$ denotes mutual information.
@@
A.3. Physics-Informed Neural ODEs
@@
Latent Space Transport Dynamics: We model the evolution of latent representations $\mathbf{x}_{i}\in\mathbb{R}^{D}$ along the latitude dimension using a transport-inspired ODE:
(26) $\frac{d\mathbf{x}_{i}}{dt}=\gamma\cdot\tanh(\mathbf{R}_{i})$
(27) $\small\mathbf{R}_{i}={\color[rgb]{0.0078125,0.46484375,0.7421875}\underbrace{\boldsymbol{\nu}\odot(\mathbf{x}_{i+1}-2\mathbf{x}_{i}+\mathbf{x}_{i-1})}_{\text{diffusion}}}+{\color[rgb]{0.84765625,0.26171875,0.08203125}\underbrace{\boldsymbol{\mu}\odot\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i-1}}{2}}_{\text{advection}}}+{\color[rgb]{0.1875,0.24609375,0.625}\underbrace{\boldsymbol{f}}_{\text{forcing}}}+{\color[rgb]{0,0.5390625,0.48046875}\underbrace{\alpha\cdot\text{MLP}(\mathbf{x}_{i})}_{\text{neural correction}}}$}
Proposition A.5 (Latent Space Transport Interpretation).
@@
(28) $\frac{d\mathbf{x}}{dt}=\mathcal{T}_{transport}(\mathbf{x};\boldsymbol{\theta}_{phys})+\mathcal{T}_{neural}(\mathbf{x};\boldsymbol{\theta}_{nn})$
where $\mathcal{T}_{transport}$ captures simplified transport processes with learnable physics parameters $\boldsymbol{\theta}_{phys}=\{\boldsymbol{\nu},\boldsymbol{\mu},\boldsymbol{f}\}\in\mathbb{R}^{3D}$, and $\mathcal{T}_{neural}$ provides additional modeling capacity through neural network parameters $\boldsymbol{\theta}_{nn}$.
@@
(29) ${\color[rgb]{0.0078125,0.46484375,0.7421875}\boldsymbol{\nu}\odot(\mathbf{x}_{i+1}-2\mathbf{x}_{i}+\mathbf{x}_{i-1})}$ $\longleftrightarrow\text{Meridional mixing and smoothing}$
${\color[rgb]{0.84765625,0.26171875,0.08203125}\boldsymbol{\mu}\odot\frac{\mathbf{x}_{i+1}-\mathbf{x}_{i-1}}{2}}$ $\longleftrightarrow\text{Meridional transport processes}$
${\color[rgb]{0.1875,0.24609375,0.625}\boldsymbol{f}}$ $\longleftrightarrow\text{External climate drivers}$
${\color[rgb]{0,0.5390625,0.48046875}\alpha\cdot\text{MLP}(\mathbf{x}_{i})}$ $\longleftrightarrow\text{Unresolved nonlinear dynamics}$
Proposition A.6 (Numerical Stability).
The transport-inspired ODE with hyperbolic tangent activation is globally Lipschitz continuous with constant $L=\gamma$, ensuring controlled error propagation:
(30) $\left\|\frac{d\mathbf{x}}{dt}-\frac{d\mathbf{x}^{\prime}}{dt}\right\|\leq\gamma\|\mathbf{x}-\mathbf{x}^{\prime}\|$
for any latent states $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{HD}$. This global Lipschitz property guarantees existence and uniqueness of solutions and supports numerical stability during integration.
Proposition A.7 (Parameter Efficiency).
The physics-inspired structure reduces the effective parameter space by constraining the dynamics to transport-consistent manifolds. Specifically, the constraint reduces the degrees of freedom from $\mathcal{O}(HD^{2})$ (unconstrained) to $\mathcal{O}(HD+D^{2})$ (transport-constrained), potentially improving sample efficiency.
Implementation Considerations: We apply periodic boundary conditions consistent with our latitude-based processing and employ adaptive step-size integration methods for numerical accuracy. The scaling factor $\gamma$ and hyperbolic tangent activation prevent gradient explosion while preserving the fundamental transport structure.
@@
A.4. Teleconnection Representation Theory
@@
Proposition A.8 (Nonlinear Teleconnection Learning).
The learned teleconnection patterns $\{\mathbf{P}_{j}\}_{j=1}^{n_{p}}\subset\mathbb{R}^{HD}$ converge to a basis that captures the dominant modes of atmospheric covariance, encompassing both linear relationships and nonlinear teleconnection patterns. Formally:
(31) $\lim_{T\to\infty}\text{span}(\{\mathbf{P}_{j}\})\supseteq\text{span}(\{\text{EOF}_{j}^{(linear)}\})$
@@
The teleconnection attention computes pattern weights as $\boldsymbol{\omega}=\text{softmax}(\mathbf{\bar{x}}\mathbf{W}^{p})$ where $\mathbf{\bar{x}}\in\mathbb{R}^{D}$ is the global mean state. During training, the patterns $\mathbf{P}_{j}$ and projection matrix $\mathbf{W}^{p}\in\mathbb{R}^{D\times n_{p}}$ are jointly optimized to minimize prediction error.
@@
Proposition A.9 (Sample Complexity for Teleconnection Learning).
To learn $k$ dominant teleconnection patterns with approximation error $\varepsilon$ in the $L_{2}$ norm, the teleconnection module requires $\mathcal{O}(k\log(D)/\varepsilon^{2})$ training samples, where the logarithmic dependence on dimension reflects the low-intrinsic-dimensional nature of atmospheric teleconnections.
Proposition A.10 (Convergence Rate).
Under standard regularity conditions on the atmospheric data distribution, the teleconnection patterns converge to their population optima at rate $\mathcal{O}(T^{-1/2})$ where $T$ is the number of training iterations, matching the convergence rate of stochastic gradient methods for smooth objectives.
@@
A.5. Complexity and Generalization Analysis
@@
A.5.1. Computational Complexity
Proposition A.11 (Time Complexity).
@@ A.5.1. Computational Complexity
A.5.2. Error Propagation Analysis
Proposition A.12 (Multi-Scale Error Bounds).
For a signal decomposed into $L+1$ frequency bands, the prediction error is bounded by:
(33) $\mathbb{E}[\|y-\hat{y}\|^{2}]\leq\sum_{\ell=0}^{L}\lambda_{\ell}\mathbb{E}[\|y_{\ell}-\hat{y}_{\ell}\|^{2}]+\mathcal{O}(L^{-\alpha})$
where $y_{\ell},\hat{y}_{\ell}$ represent ground truth and prediction for frequency band $\ell$, $\lambda_{\ell}\geq 0$ are band-specific weights with $\sum_{\ell}\lambda_{\ell}=1$, and the additional term bounds decomposition error.
@@ A.5.2. Error Propagation Analysis
The multi-scale architecture processes frequency bands independently before fusion, creating a natural error decomposition. The hierarchical fusion mechanism that privileges low-frequency information creates weights $\lambda_{\ell}$ that typically decrease with frequency, providing robustness against high-frequency noise. The decomposition error term reflects the finite precision of the learned wavelet basis. ∎
A.5.3. Generalization Bounds
Proposition A.13 (Generalization Error).
For a training set of size $n$ drawn from the atmospheric state distribution, with probability at least $1-\delta$, the generalization error satisfies:
(34) $\mathcal{R}(h)-\hat{\mathcal{R}}_{n}(h)\leq\mathcal{O}\left(\sqrt{\frac{\log(L+1)+\sum_{\ell=0}^{L}\mathcal{C}(h_{\ell})+\log(1/\delta)}{n}}\right)$
where $\mathcal{R}(h)$ is the true risk, $\hat{\mathcal{R}}_{n}(h)$ is the empirical risk, and $\mathcal{C}(h_{\ell})$ measures the complexity of the model component for frequency band $\ell$.